OpenStudy (anonymous):

f(x)=(x-5)\(x+5) Find the inverse.

OpenStudy (anonymous):

to find the inverse of a function, dont you switch all the Ys and Xs and then solve for Y?

OpenStudy (anonymous):

yes but i dont know how to solve (y-5)/(y+5)

OpenStudy (anonymous):

-5(x+1)/(x-1)

OpenStudy (anonymous):

im pretty sure you need to solve for x first then switch Xs with Ys.. but i still dont know how to do that

OpenStudy (anonymous):

y = (x-5)/(x+5) x = (y-5)/(y+5) x (y+5) = y - 5 xy + 5x = y - 5 xy - y = -5x - 5 y (x - 1) = -5(x+1) y = -5(x+1)/(x-1)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so i just cross multiply

OpenStudy (anonymous):

indeed, cross multiply, distribute out all the terms, then collect the y terms on one side. factor out a y, and bring all the x's back to the other side. it's kinda like... start out with everything separate, mash it all together, then separate it all out again

OpenStudy (anonymous):

got it.. u da man